# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v2

@@ Abstract
Bacteria evolve in volatile environments and complex spatial structures. Migration, fluctuations and environmental variability therefore have a significant impact on the evolution of microbial populations. Here, we consider a class of spatially explicit metapopulation models arranged as regular (circulation) graphs where wild-type and mutant cells compete in a time-fluctuating environment where demes (subpopulations) are connected by slow cell migration. The carrying capacity is the same at each deme and endlessly switches between two values associated to harsh and mild environmental conditions. It is known that environmental variability can lead to population bottlenecks, following which the population is prone to fluctuation-induced extinction. Here, we analyse how slow migration, spatial structure, and fluctuations affect the phenomena of fixation and extinction on clique, cycle, and square lattice metapopulations. When the carrying capacity remains large, bottlenecks are weak and deme extinction can be ignored. The dynamics is thus captured by a coarse-grained description within which the probability and mean time of fixation are obtained analytically. This allows us to show that, in contrast to what happens in static environments, the mutant fixation probability depends on the rate of migration. We also show that the fixation probability and mean fixation time can exhibit a non-monotonic dependence on the switching rate. When the carrying capacity is small under harsh conditions, bottlenecks are strong, and the metapopulation evolution is shaped by the coupling of deme extinction and strain competition. This yields rich dynamical scenarios, among which we identify the best conditions to eradicate mutants without dooming the metapopulation to extinction. We offer an interpretation of these findings in the context of an idealised treatment strategy and discuss possible generalisations of our models.
@@ II Model & Methods
Figure 1: Metapopulation dynamics in a static environment. (a) Examples of metapopulation graphs: a clique, cycle, and grid (from left to right). Neighbouring demes are connected by migration (double arrows). Initially, there is one mutant deme (red/light) and $\Omega-1$ wild-type demes (blue / dark), and all demes have the same constant carrying capacity $K$. (b) Dynamics in a single deme. Left: Wild-type $W$ cells (blue / dark) compete with mutants of type $M$ (red / light). When $K$ is small, the deme is prone to extinction. When $K$ is large, both types coexist prior to $W$ or $M$ fixation. Top right: Realisations of the rescaled deme size $n/K$ vs. time $t$ for $K=5$ (orange/light) and $K=100$ (green/dark) illustrating how $n$ fluctuates about $K$. Bottom right: Fraction of $M$ cells vs. $t$ in a deme with $K=100$. Deme extinction is not observed. Transient coexistence of $W$ and $M$ is followed by the fixation of $W$ (blue traces) or $M$ (red traces). Here $s=0.01$. (c) Invasion of $W$ deme by an $M$ cell: Any $M$ cell of a deme migrates to a neighbouring $W$ site with migration rate $m$ after a mean time $\Delta t=1/(mK)$, and then type $M$ either quickly fixates, producing a new $M$ deme (right), or does not fix leaving the pair of $M$ and $W$ demes unchanged (left). The same picture holds for the invasion of an $M$ deme by a $W$ cell; see text. (d) Intermediate regime (here for the clique): The dynamics is characterised by deme extinction and $W/M$ competition (see text). Deme extinction occurs after a mean time $\tau_{E}$, and empty demes are then recolonised by an invader from a neighbouring surviving deme after $\Delta t\sim 1/(mK)$. A recolonised deme is rapidly taken over (in $\Delta t\sim\mathcal{O}(\ln(K))$). (e) Coarse-grained description of the metapopulation dynamics: Each deme is always either fully $W$ (blue / dark) or $M$ (red / light) or empty (white). In this description, different scenarios arise, shown for the clique. Competition-dominated regime: all demes are occupied and there is always fixation of $W$ or $M$. Intermediate regime: eventually $W$ or $M$ take over all occupied demes, resulting in a dynamic equilibrium of empty demes and $W$ or $M$ demes; see text. Extinction-dominated regime: there are frequent deme extinctions and the metapopulation quickly goes extinct.
@@
The microbial metapopulation can be envisioned as a graph whose nodes $x\in\{1,\dots,\Omega\}$ are demes (also called sites). Each deme $x$ is a well-mixed subpopulation of size $n(x)=n_{W}(x)+n_{M}(x)$ located at a node of the metapopulation graph. Thus, $N_{W/M}\equiv\sum_{x}n_{W/M}(x)$ is the number of $W/M$ individuals across the metapopulation, and $N\equiv N_{W}+N_{M}$ is the total number of individuals in the whole graph. Here, we focus on fully-connected graphs (as in the island model  [33, 34]), called cliques, and periodic one- and two-dimensional lattices called respectively cycles and grids; see Fig. 1(a). These are regular circulation graphs, generally denoted by ${\rm G}=\{{\rm clique,cycle,grid}\}$, of $\Omega$ demes connected by edges to their $q_{{\rm G}}$ nearest neighbours via cell migration at per capita rate $m$ (independently from division and death) [33, 34, 43, 41, 42, 60]; see Fig. 1(a-c). We study the eco-evolutionary dynamics of the metapopulation in the biologically relevant regime of slow migration (whereby intra-deme dynamics occur much faster than inter-deme dynamics; see below) [61, 62, 41, 42, 63], and consider that initially one deme is occupied entirely by mutants ($M$ deme), while the other $\Omega-1$ demes ($W$ demes) are all populated by $W$ cells; see Sec. V. All demes are assumed to have the same carrying capacity $K$ which encodes environmental variability. In Sec. III, we assume that $K$ is constant, and in Sec. IV we let the carrying capacity switch endlessly between two values representing mild and harsh conditions [16, 17, 18, 23, 20, 19, 21, 22, 24, 32]; see below.
@@
The inter-deme dynamics stems from the migration of one cell of type $\alpha\in\{W,M\}$ from the site $x$ to one of its $q_{\rm G}$ neighbouring demes denoted by $y$ at a per-capita migration rate $m$. Here, for the sake of simplicity, we assume that the migration rate is the same in all directions and for both types (symmetric migration) (see Sec. V for a discussion of these assumptions). The inter-deme dynamics for all cells at deme $x$ with its neighbouring demes labelled $y$ is therefore implemented according to the reaction
@@
With this, the condition for a metapopulation to be a circulation is given by
@@
$K(t)=\frac{1}{2}\left[K_{-}+K_{-}+\xi(t)(K_{+}-K_{-})\right].$ (6)
@@ III Static environments
when $K\gg 1$; see Fig. 2(e). We thus refer to $K$ as “small” where fixation of a deme (intra-deme dynamics) occurs on slower timescale than its extinction. Similarly, $K$ is referred to as large (i.e. $K\gg 1$) where fixation of a site occurs faster than deme extinction. In our analysis, we distinguish between different dynamical scenarios through
@@
The expression of $\Omega_{\text{occ}}$ ignores spatial correlations and hence is not accurate for cycles and grids if $\psi(m,K)$ is not much larger than 1. However, $\psi(m,K)$ allows us to efficiently distinguish between the regimes dominated by $M/W$ competition, deme extinction, and the crossover intermediate regime; see Fig. 9 and Appendix D.
Henceforth, we refer to “invasion” when a cell of type $M/W$ migrates to and fixates in a $W/M$ deme, and to “recolonisation” when a cell of either type migrates into an empty deme and takes it over; see Fig. 1(c,d). Accordingly, the competition-dominated and intermediate regimes are respectively characterised by invasions and recolonisations; see Fig. 1(c,d). A summary of the timescales for the different processes and regimes considered in this work can be found in Tables 1 and 2.
@@ III.1 Competition-dominated dynamics
Figure 2: (a-d): Competition-dominated dynamics, $\psi\gg 1$. (a) $M$ fixation probability $\phi$ vs. constant carrying capacity $K$; (b) unconditional mean fixation time $\theta$ vs. $K$; (c) $\phi$ vs. per capita migration rate $m$; (d) $\theta$ vs. $m$. Markers are simulation results and lines are predictions of Eq. (14) for $s=0.1$ (blue) and $s=0.01$ (red) on a clique (solid lines / crosses), cycle (dashed lines / circles), and grid (dotted lines / triangles). In (a,b), $m=10^{-4},\Omega=16$, and in (c,d), $K=50,\Omega=16$. In (a,c), markers for the same $s$ are almost indistinguishable indicating independence of the spatial structure. (e,f): Extinction-dominated dynamics, $\psi<1$. (e) Mean extinction time of a single deme $\tau_{E}$ vs. $K$ ($m=0$). Circles are simulation data, line shows the predictions of Eq. (7). (f) Metapopulation mean extinction time $\theta_{E}$ vs. $K$ for $\Omega=16$ and $m=10^{-2}$ (blue) and $10^{-4}$ (red). Markers are simulation results and thick lines are predictions of Eq. (39) for cliques (solid lines / crosses), cycles (dashed lines / circles), and grids (dotted lines / triangles). Thin dashed vertical lines are guides to the eye showing $\psi=1$ for $m=10^{-2}$ (blue) and $10^{-4}$ (red). Selection plays no role in this regime, so panels (e,f) have been obtained with $s=0$. In all panels, there is initially one $M$ deme and $\Omega-1$ demes occupied by $W$. In panels (a,c,f), dashed lines overlap with solid lines and so are not visible. Error bars are plotted in each case but are typically too small to see.
When $\psi\gg 1$ with $m\ll 1$ (slow migration; see below), the carrying capacity is large enough for many migrations to occur on the timescale $\tau_{E}$ of deme extinction, with intra-deme dynamics occurring faster than inter-deme dynamics. Since every deme expects many incoming cells in time $\tau_{E}$, deme extinction is unlikely and can be neglected. In this regime, the dynamics is dominated by local $M/W$ competition: $W$ and $M$ cells compete in each deme to fixate the local subpopulation; see Fig. 1(b).
As in Refs. [41, 51, 42], the slow migration condition allows us to adopt a coarse-grained description treating each deme as a single entity of type $W$ or $M$. This is because in the regime of slow migration, the mean time for an $M$ or $W$ invader to fixate a deme is much shorter compared to $1/(mK)$, the expected time between migrations (see Appendix C, Table 1 and Eq. 12). When the dynamics is dominated by $W/M$ competition and $m\ll 1$, at coarse-grained level, each deme can thus be regarded as being either entirely occupied by $W$ or $M$ individuals; see Fig. 1(c,e). In this regime, each sequential migration is an invasion attempt, with a cell from an $M/W$ site trying to invade a neighbouring $W/M$ deme; see Fig. 1(c). Here, an $M/W$ invasion is the fixation of a single $M/W$ mutant in a deme consisting of K1K-1 cells of type $W/M$.
@@
These rates, together with the fact that the timescale of intra-deme dynamics is $1/s$ (timescale of fixation of a single isolated deme [64, 7, 65, 66, 67]; see Table 1) can be used to define slow migration in the competition-dominated regime. To this end, we notice that the growth of the cluster of $i$ mutant demes in time $1/s$ is $T^{+}_{i}(m,{\rm G},K)/s=mKE_{{\rm G}}(i)\rho_{M}/(q_{{\rm G}}s)$. [^1] Here, we use the growth of the $M$ cluster to define slow migration as the mutant has a fitness advantage. Therefore, $\rho_{M}>\rho_{W}$ and $T^{+}_{i}(m,\mathrm{G},K)>T_{i}^{-}(m,\mathrm{G},K)$ in general. In the coarse-grained description, slow migration is the regime where invasion can be regarded as being instantaneous, with fixation of a successful $M$ invader occurring before the next invasion. This requires that the average number of successful $M$ invaders in the time for the intra-deme dynamics be less than one, i.e. $T^{+}_{i}(m,{\rm G},K)/s<1$. Therefore, the condition for slow migration here is
@@
where we have used $q_{\rm G}/E_{\rm G}(i)\leq 1$. When $s\ll 1$ and $Ks\gg 1$, we have $\rho_{M}\sim s$ [64, 7, 65, 66, 67] and there is slow migration when $m\lesssim 1/K$. For typical values used here, e.g. $\Omega=16,K=100$ and $s=0.1$, we estimate that there is slow migration if $m\lesssim 10^{-2}$, which is in line with the values of $m\in[10^{-5},10^{-2}]$ used in our examples, and explains the deviations reported in Fig. 2(d) for larger $m$. The coarse-grained competition-dominated dynamics is thus a birth-death process for the number $i$ of $M$ demes, with absorbing boundaries at $i=\Omega$ ($M$ fixation) and $i=0$ ($W$ fixation); see Appendix G. In this representation, the $M$ fixation probability in a metapopulation of size $\Omega$, spatially structured as a graph $\rm G$}, consisting initially of $i$ mutant demes is denoted $\phi_{i}^{{\rm G}}$, and the unconditional (i.e. regardless of whether $M$ or $W$ takes over [7, 67, 66]) mean fixation time (uMFT) denoted $\theta_{i}^{{\rm G}}$. These quantities satisfy the first-step equations [72, 7, 67, 66]
@@
where $\gamma\equiv T_{i}^{-}/T_{i}^{+}=\rho_{W}/\rho_{M}\approx\exp(-Ks)$ is a quantity independent of $m$ and $\rm G$}. As noted in Refs. [41, 50, 42] the fixation probability $\phi^{{\rm G}}=\phi$ is therefore independent of the migration rate and spatial structure. This remarkable result stems from the graphs considered here being circulations; see Eq. (5). In static environments, a generalised circulation theorem ensures that the fixation probability is independent of $m$ and $\rm G$} for circulation graphs [47, 41, 42], a feature displayed in the stochastic simulations of Fig. 2(a,c) for the full microscopic model. In excellent agreement with simulation data of Fig. 2(a,c), we find that the $M$ fixation probability increases almost exponentially with $Ks$ and approaches 11 when $Ks\gg 1$, $\phi^{{\rm G}}\approx 1$. This stems from the invasion of $W$ demes being increasingly likely (and the invasion of $M$ demes exponentially less likely) when the average number of migrations ($mK$) increases along with $K$. When $Ks\ll 1$, the competition is effectively neutral, and in this case $\phi^{{\rm G}}\approx 1/\Omega$. In good agreement with simulation results of Fig. 2(b,d), Eq. (14) predicts that the uMFT decreases with the migration rate $\theta^{{\rm G}}\sim 1/m$ and, for given parameters, the uMFT is shortest on cliques, while it is larger on cycles than on grids. Intuitively, for higher $m$ and more connected graphs, migrants spread faster leading to quicker invasion and fixation.

### v3

@@ Abstract
Bacteria evolve in volatile environments and complex spatial structures. Migration, fluctuations and environmental variability therefore have a significant impact on the evolution of microbial populations. Here, we consider a class of spatially explicit metapopulation models arranged as regular (circulation) graphs where wild-type and mutant cells compete in a time-fluctuating environment in which demes (subpopulations) are connected by slow cell migration. The carrying capacity is the same at each deme and endlessly switches between two values associated to harsh and mild environmental conditions. It is known that environmental variability can lead to population bottlenecks, following which the population is prone to fluctuation-induced extinction. Here, we analyse how slow migration, spatial structure, and fluctuations affect the phenomena of fixation and extinction on clique, cycle, and square lattice metapopulations. When the carrying capacity remains large, bottlenecks are weak and deme extinction can be ignored. The dynamics is thus captured by a coarse-grained description within which the probability and mean time of fixation are obtained analytically. This allows us to show that, in contrast to what happens in static environments, the mutant fixation probability depends on the rate of migration. We also show that the fixation probability and mean fixation time can exhibit a non-monotonic dependence on the switching rate. When the carrying capacity is small under harsh conditions, bottlenecks are strong, and the metapopulation evolution is shaped by the coupling of deme extinction and strain competition. This yields rich dynamical scenarios, among which we identify the best conditions to eradicate mutants without dooming the metapopulation to extinction. We offer an interpretation of these findings in the context of an idealised treatment strategy and discuss possible generalisations of our models.
@@ II Model & Methods
Figure 1: Metapopulation dynamics in a static environment. (a) Examples of metapopulation graphs: a clique, cycle, and grid (from left to right). Neighbouring demes are connected by migration (double arrows). Initially, there is one mutant deme (red/light) and $\Omega-1$ wild-type demes (blue / dark), and all demes have the same constant carrying capacity $K$. (b) Dynamics in a single deme. Left: Wild-type $W$ cells (blue / dark) compete with mutants of type $M$ (red / light). When $K$ is small, the deme is prone to extinction. When $K$ is large, both types coexist prior to $W$ or $M$ fixation. Top right: Realisations of the rescaled deme size $n/K$ vs. time $t$ for $K=5$ (orange/light) and $K=100$ (green/dark) illustrating how $n$ fluctuates about $K$. Bottom right: Fraction of $M$ cells vs. $t$ in a deme with $K=100$. Deme extinction is not observed. Transient coexistence of $W$ and $M$ is followed by the fixation of $W$ (blue traces) or $M$ (red traces). Here $s=0.01$. (c) Invasion of $W$ deme by an $M$ cell: Any $M$ cell of a deme migrates to a neighbouring $W$ site with migration rate $m$ after a mean time $\Delta t=1/(mK)$, and then type $M$ either quickly fixates, producing a new $M$ deme (right), or does not fix leaving the pair of $M$ and $W$ demes unchanged (left). The same picture holds for the invasion of an $M$ deme by a $W$ cell; see text. (d) Deme recolonisation (here for the clique): Deme extinction occurs after a mean time $\tau_{E}$, and empty demes are then recolonised by an invader from a neighbouring surviving deme after $\Delta t\sim 1/(mK)$. A recolonised deme is rapidly taken over (in $\Delta t\sim\mathcal{O}(\ln(K))$). (e) Coarse-grained description of the metapopulation dynamics: Each deme is always either fully $W$ (blue / dark) or $M$ (red / light) or empty (white). In this description, different scenarios arise, shown for the clique. Competition-dominated regime: all demes are occupied and there is always fixation of $W$ or $M$. Extinction-dominated regime: there are frequent deme extinctions and the metapopulation quickly goes extinct.
@@
The microbial metapopulation can be envisioned as a graph whose nodes $x\in\{1,\dots,\Omega\}$ are demes (also called sites). Each deme $x$ is a well-mixed subpopulation of size $n(x)=\sum_{\alpha\in\{W,M\}}n_{\alpha}(x)=n_{W}(x)+n_{M}(x)$ located at a node of the metapopulation graph. Thus, $N_{W/M}\equiv\sum_{x}n_{W/M}(x)$ is the number of $W/M$ individuals across the metapopulation, and $N\equiv N_{W}+N_{M}$ is the total number of individuals in the whole graph. Here, we focus on fully-connected graphs (as in the island model  [33, 34]), called cliques, and periodic one- and two-dimensional lattices called respectively cycles and grids; see Fig. 1(a). These are regular graphs, generally denoted by ${\rm G}=\{{\rm clique,cycle,grid}\}$, of $\Omega$ demes connected by edges to their $q_{{\rm G}}$ nearest neighbours via cell migration at per capita rate $m$ (independently from division and death) [33, 34, 43, 41, 42, 60]; see Fig. 1(a-c). We study the eco-evolutionary dynamics of the metapopulation in the biologically relevant regime of slow migration (whereby intra-deme dynamics occur much faster than inter-deme dynamics; see below) [61, 62, 41, 42, 63], and consider that initially one deme is occupied entirely by mutants ($M$ deme), while the other $\Omega-1$ demes ($W$ demes) are all populated by $W$ cells; see Sec. V. All demes are assumed to have the same carrying capacity $K$ which encodes environmental variability. In Sec. III, we assume that $K$ is constant, and in Sec. IV we let the carrying capacity switch endlessly between two values representing mild and harsh conditions [16, 17, 18, 23, 20, 19, 21, 22, 24, 32]; see below.
@@
The inter-deme dynamics stems from the migration of one cell of type $\alpha\in\{W,M\}$ from the site $x$ to one of its $q_{\rm G}$ neighbouring demes denoted by $y$ at a per-capita migration rate $m$. Here, for the sake of simplicity, we assume that the migration rate is the same in all directions and for both types (symmetric migration); see Sec. V for a discussion of these assumptions. The inter-deme dynamics for all cells at deme $x$ with its neighbouring demes labelled $y$ is therefore implemented according to the reaction
@@
On a given metapopulation, if the number of cells migrating into a deme equals the number of cells migrating out, and this is true for all demes, we say that the metapopulation is a circulation. Precisely, the condition here for a metapopulation to be a circulation is given by
@@
$K(t)=\frac{1}{2}\left[K_{-}+K_{-}+\xi(t)(K_{+}-K_{-})\right],$ (6)
@@ III Static environments
when $K\gg 1$; see Fig. 2(b,top). We thus refer to $K$ as “small” where fixation of a deme (intra-deme dynamics) occurs on slower timescale than its extinction. Similarly, $K$ is referred to as large (i.e. $K\gg 1$) where fixation of a site occurs faster than deme extinction. In our analysis, we distinguish between different dynamical scenarios through
@@
The expression of $\Omega_{\text{occ}}$ ignores spatial correlations and hence is not accurate for cycles and grids if $\psi(m,K)$ is not much larger than 1. However, $\psi(m,K)$ allows us to efficiently distinguish between the regimes dominated by $M/W$ competition ($\psi\gg 1$) and deme extinction ($\psi<1$); see Appendix D. The crossover intermediate regime ($\psi\gtrsim 1$) is discussed in detail in Appendix F.
Henceforth, we refer to “invasion” when a cell of type $M/W$ migrates to and fixates in a $W/M$ deme, and to “recolonisation” when a cell of either type migrates into an empty deme and takes it over; see Fig. 1(c,d). The competition-dominated dynamics is characterised by invasions, while the extinction-dominated regime consists of extinctions and some recolonisations prior to the final extinction of the metapopulation. A summary of the timescales for the different processes and regimes considered in this work can be found in Tables 1 and 2.
@@ III.1 Competition-dominated dynamics
Figure 2: (a): Competition-dominated dynamics, $\psi\gg 1$. (Top left) $M$ fixation probability $\phi$ vs. constant carrying capacity $K$; (Bottom left) unconditional mean fixation time $\theta$ vs. $K$; (Top right) $\phi$ vs. per capita migration rate $m$; (Bottom right) $\theta$ vs. $m$. Markers are simulation results and lines are predictions of Eq. (14) for $s=0.1$ (blue) and $s=0.01$ (red) on a clique (solid lines / crosses), cycle (dashed lines / circles), and grid (dotted lines / triangles). In (left), $m=10^{-4},\Omega=16$, and in (right), $K=50,\Omega=16$. In (top), markers for the same $s$ are almost indistinguishable indicating independence of the spatial structure. (b): Extinction-dominated dynamics, $\psi<1$. (Top) Mean extinction time of a single deme $\tau_{E}$ vs. $K$ ($m=0$). Circles are simulation data, line shows the predictions of Eq. (7). (Bottom) Metapopulation mean extinction time $\theta_{E}$ vs. $K$ for $\Omega=16$ and $m=10^{-2}$ (blue) and $10^{-4}$ (red). Markers are simulation results and thick lines are predictions of Eq. (39) for cliques (solid lines / crosses), cycles (dashed lines / circles), and grids (dotted lines / triangles). Thin dashed vertical lines are guides to the eye showing $\psi=1$ for $m=10^{-2}$ (blue) and $10^{-4}$ (red). Selection plays no role in this regime, so simulation data for (b) has been obtained with $s=0$. In all panels, there is initially one $M$ deme and $\Omega-1$ demes occupied by $W$. In panels (a,top) and (b,bottom), dashed lines overlap with solid lines and so are not visible. Error bars are plotted in each case but are typically too small to see.
When $\psi\gg 1$ with $mK<1$ (slow migration; see Eq. (12)), the carrying capacity is large enough for many migrations to occur on the timescale $\tau_{E}$ of deme extinction, with intra-deme dynamics occurring faster than inter-deme dynamics. Since every deme expects many incoming cells in time $\tau_{E}$, deme extinction is unlikely and can be neglected. In this regime, the dynamics is dominated by local $M/W$ competition: $W$ and $M$ cells compete in each deme to fixate the local subpopulation; see Fig. 1(b).
As in Refs. [41, 51, 42], the slow migration condition allows us to adopt a coarse-grained description treating each deme as a single entity of type $W$ or $M$. This is because in the regime of slow migration, the mean time for an $M$ or $W$ invader to fixate a deme is much shorter compared to $1/(mK)$, the expected time between migrations (see Appendix C, Table 1 and Eq. 12). When the dynamics is dominated by $W/M$ competition and $mK<1$, at a coarse-grained level, each deme can thus be regarded as being either entirely occupied by $W$ or $M$ individuals; see Fig. 1(c,e). In this regime, each sequential migration is an invasion attempt, with a cell from an $M/W$ site trying to invade a neighbouring $W/M$ deme; see Fig. 1(c). Here, an $M/W$ invasion is the fixation of a single $M/W$ mutant in a deme consisting of K1K-1 cells of type $W/M$.
@@
These rates, together with the fact that the timescale of intra-deme dynamics is $1/s$ (timescale of fixation of a single isolated deme [64, 7, 65, 66, 67]; see Table 1) can be used to define slow migration in the competition-dominated regime. To this end, we notice that the growth of the cluster of $i$ mutant demes in time $1/s$ is $T^{+}_{i}(m,{\rm G},K)/s=mKE_{{\rm G}}(i)\rho_{M}/(q_{{\rm G}}s)$. [^1] Here, we use the growth of the $M$ cluster to define slow migration as the mutant has a fitness advantage. Therefore, $\rho_{M}>\rho_{W}$ and $T^{+}_{i}(m,\mathrm{G},K)>T_{i}^{-}(m,\mathrm{G},K)$ in general. In the coarse-grained description, slow migration is the regime where invasion can be regarded as being instantaneous, with fixation of a successful $M$ invader occurring before the next invasion. This requires that the average number of successful $M$ invasions on the timescale of the intra-deme dynamics is less than one, i.e. $T^{+}_{i}(m,{\rm G},K)/s<1$. Therefore, the condition for slow migration here is
@@
where we have used $q_{\rm G}/E_{\rm G}(i)\leq 1$. When $s\ll 1$ and $Ks\gg 1$, we have $\rho_{M}\sim s$ [64, 7, 65, 66, 67] and there is slow migration when $mK\lesssim 1$, i.e. the expected number of migrants from a deme per unit time is less than one. For typical values used here, e.g. $\Omega=16,K=100$ and $s=0.1$, we estimate that there is slow migration if $m\lesssim 10^{-2}$, which is in line with the values of $m\in[10^{-5},10^{-2}]$ used in our examples, and explains the deviations reported in Fig. 2(a,bottom right) for larger $m$. The coarse-grained competition-dominated dynamics is thus a birth-death process for the number $i$ of $M$ demes, with absorbing boundaries at $i=\Omega$ ($M$ fixation) and $i=0$ ($W$ fixation); see Appendix G. In this representation, the $M$ fixation probability in a metapopulation of size $\Omega$, spatially structured as a graph $\rm G$}, consisting initially of $i$ mutant demes is denoted $\phi_{i}^{{\rm G}}$, and the unconditional (i.e. regardless of whether $M$ or $W$ takes over [7, 67, 66]) mean fixation time (uMFT) denoted $\theta_{i}^{{\rm G}}$. These quantities satisfy the first-step equations [72, 7, 67, 66]
@@
where $\gamma\equiv T_{i}^{-}/T_{i}^{+}=\rho_{W}/\rho_{M}\approx\exp(-Ks)$ is a quantity independent of $m$ and $\rm G$}. As noted in Refs. [41, 50, 42] the fixation probability $\phi^{{\rm G}}=\phi$ is therefore independent of the migration rate and spatial structure. This remarkable result stems from the graphs considered here being circulations; see Eq. (5). In static environments, a generalised circulation theorem ensures that the fixation probability is independent of $m$ and $\rm G$} for circulation graphs [47, 41, 42], a feature displayed in the stochastic simulations of Fig. 2(a,top) for the full microscopic model. In excellent agreement with simulation data of Fig. 2(a,top), we find that the $M$ fixation probability increases almost exponentially with $Ks$ and approaches 11 when $Ks\gg 1$, $\phi^{{\rm G}}\approx 1$. This stems from the invasion of $W$ demes being increasingly likely (and the invasion of $M$ demes exponentially less likely) when the average number of migrations ($mK$) increases along with $K$. When $Ks\ll 1$, the competition is effectively neutral, and in this case $\phi^{{\rm G}}\approx 1/\Omega$. In good agreement with simulation results of Fig. 2(a,bottom), Eq. (14) predicts that the uMFT decreases with the migration rate $\theta^{{\rm G}}\sim 1/m$ and, for given parameters, the uMFT is shortest on cliques, while it is larger on cycles than on grids. Intuitively, for higher $m$ and more connected graphs, migrants spread faster leading to quicker invasion and fixation.